# Affiliations withheld for double-blind review

### v2

@@ Discussion
In this study, we proposed a KG-guided chain-of-thought generation framework that couples biomedical knowledge from PrimeKG with LLM-based reasoning over structured EHR data. By aligning ICD–9 codes to PrimeKG,chandak2023primekg mining disease-specific relevance sets and paths, and using these as scaffolds for CoT generation and filtering, we construct a compact supervision corpus for data-efficient fine-tuning of lightweight LLMs. On MIMIC-III visit-level disease prediction, KG-guided CoT improves both LLaMA-3.1-8B and Gemma-7B over their untuned counterparts and makes them competitive with strong classical baselines. With only 400 labeled visits, KG-CoT–tuned models reach AUROC $\approx 0.67$ and macro-AUPR $\approx 0.40$; with 1,000 visits, performance increases to AUROC $\approx 0.70$ and macro-AUPR $\approx 0.47$. These results suggest that KG-anchored CoT supervision acts as an effective reasoning prior, helping small LLMs learn clinically meaningful decision boundaries from limited data. Classical models remain strong on AUROC, but KG-guided LLMs close or surpass this gap while improving macro-AUPR. This points toward hybrid designs in which classical or deep risk models provide scores, and KG-guided LLMs act as a clinician-facing layer that produces explanations. We also observe encouraging cross-cohort generalization. In zero-shot transfer from MIMIC-III to CRADLE, fine-tuned LLaMA and Gemma models increase accuracy from roughly 0.40–0.51 to 0.72–0.77 and macro-F1 from 0.31–0.34 to 0.46–0.51, while AUROC and AUPR remain stable. This pattern indicates that discrimination is preserved, while the induced decision boundaries and reasoning adapt to the target cohort.
Human evaluation provides complementary evidence that KG guidance improves the clinical usability of explanations. In 115 blinded case comparisons between our KG-guided LLaMA model and an untuned LLM baseline, the KG-guided traces were preferred on clarity and coherence, coverage and relevance, and correctness and soundness in the large majority of cases. Qualitative feedback from the MD professor and PhD student indicated that KG-guided traces more often (i) organized evidence in a temporally consistent way, (ii) highlighted clinically meaningful risk factors rather than superficial code patterns, and (iii) avoided unsupported leaps or hallucinated details. By contrast, the untuned baseline was described as more “trigger-happy” in predicting positive outcomes (more likely to conclude YES) and sometimes reached correct labels via illogical or poorly justified reasoning, especially on shorter or noisier inputs. These observations suggest that anchoring CoT to KG paths, combined with label-consistency filtering, shapes both the style and the content of reasoning toward disease-specific, clinically grounded narratives.
Several limitations remain. Our approach depends on the coverage and quality of PrimeKG, so gaps or biases can propagate into CoT supervision. The mapping of ICD–9 to KG and the selection of the relevance/path rely on GPT-4o and heuristic thresholds, which may introduce subtle errors during data generation.
Looking ahead, our results suggest several concrete directions. Extending the framework to richer EHR views (notes, medications, labs) and additional KGs could support more nuanced multimodal reasoning, while moving beyond simple shortest paths to richer graph reasoning may improve explanation quality. In addition, clinician feedback highlights the value of better input structuring—making exposure and prediction windows explicit, distinguishing inpatient from outpatient encounters and medications, and adding basic demographics features (e.g. age, sex) —to reduce hallucinations and over-interpretation. Future work can introduce temporal or guideline-based constraints to further suppress illogical reasoning, discordant conclusions, and clinically implausible trajectories in the generated CoT.
@@ Acknowledgements
This research was partially supported by internal funds and GPU servers provided by the Computer Science Department of Emory University.
Implementation Details.
All fine-tuning experiments were conducted on a single NVIDIA H200 GPU with 512 GB host memory on Emory’s high-performance computing cluster. We fine-tuned the LLaMA-3.1-8B-Instruct and Gemma-7B-Instruct backbones on the KG-guided CoT corpus using full-parameter AdamW (learning rate $1\times 10^{-5}$, cosine schedule, max gradient norm 1.0), a per-device batch size of 1 with gradient accumulation of 8 (effective batch size 8), sequence length 4,096, and 10 training epochs in bf16 under DeepSpeed ZeRO-3. We monitored validation performance and selected the checkpoint with the best macro-AUPR. KG-guided CoT traces were generated using GPT-4o; across entity-mapping validation, node relevance selection, path pruning, and CoT generation, the total API cost for data creation was approximately \$100. The exact prompt templates used in our pipeline are listed below.
Figure 4: Prompt templates used in our KG-guided CoT pipeline: (top-left) disease-relevant node selection, (top-right) KG path pruning and selection, (bottom) CoT generation conditioned on KG evidence and visit features.

### v3

@@ Discussion
We introduce a KG-guided chain-of-thought (CoT) framework that uses PrimeKG structure to scaffold LLM reasoning over structured EHRs. By mapping ICD–9 codes to PrimeKG,2 mining disease-specific relevance sets and short paths, and applying label-consistency filtering, we build a compact supervision corpus for data-efficient fine-tuning of lightweight LLMs.
Finding 1: KG-anchored CoT improves low-data learning, especially under imbalance.
Across both 400 and 1,000 labeled index visits (Table Experiment:), KG-guided CoT substantially improves LLaMA-3.1-8B and Gemma-7B and is competitive with strong classical baselines. Gains are largest on macro-AUPR (e.g., $\approx 0.25\rightarrow 0.40$ for Gemma at 400), suggesting that KG-anchored supervision encourages disease-relevant evidence over superficial code co-occurrence.
Finding 2: Filtering and KG evidence play complementary roles.
In the ablation (Table 3), filtering alone yields the best accuracy/F1, while adding KG evidence yields the best AUROC (up to $\approx 0.70$). Practically, filtering reduces noisy supervision and spurious positives, whereas KG structure strengthens ranking; threshold calibration can further improve accuracy/F1. Future work can further improve threshold-dependent metrics (accuracy/F1) by calibrating decision thresholds per disease or cohort.
Finding 3: Zero-shot transfer mainly shifts the operating point, motivating calibration.
On CRADLE (Table 4), fine-tuned models boost accuracy (0.72–0.77) with only modest AUROC/AUPR changes, indicating a changed decision rule under a fixed threshold rather than large ranking gains. Deployment evaluations should therefore include calibration and threshold selection, especially across cohorts with different prevalences and coding practices.
Finding 4: Clinicians preferred KG-guided explanations.
Clinicians preferred KG-guided traces on clarity/coherence, coverage/relevance, and correctness/soundness for most cases. They cited better temporal organization, more clinically meaningful risk factors, and fewer unsupported leaps, suggesting KG evidence as an effective planning scaffold for clinician-facing rationales.
Limitations and future work.
Our approach depends on the coverage and quality of PrimeKG, so gaps or biases in the KG can propagate into both supervision and explanations. The ICD to KG mapping and the relevance/path selection steps rely on GPT-4o and heuristic constraints. While we constrain candidate paths using shortest-path enumeration and use the LLM only as a selector, the pruned set is not guaranteed to be exhaustive: alternative clinically valid paths may be removed in favor of a compact scaffold. Methodologically, we also adopt a single-visit formulation ($x_{t}$) for next-visit prediction, which simplifies evaluation but does not fully exploit longitudinal history. Future work can extend the framework to richer EHR views (notes, medications, labs), incorporate multi-visit context, and integrate clinician-in-the-loop validation for KG evidence selection and pruning.
@@ Acknowledgements
This research was partially supported by the internal funds and GPU servers provided by the Computer Science Department of Emory University, the US National Science Foundation under Award Numbers 2442172, 2312502, 2319449, and the US National Institutes of Health under Award Numbers K25DK135913, RF1NS139325, R01DK143456 and U18DP006922.